# Stochastic Optimization with Random Search

### v1

@@ 1 Introduction
Zeroth-order (derivative-free) optimization refers to a class of algorithms that minimize an objective function using only function evaluations. These methods are crucial in modern machine learning applications, where gradients are either inaccessible, unreliable, or too costly to obtain. Such scenarios arise in black-box adversarial attacks [Chen et al., 2017, Ughi et al., 2022], hyperparameter tuning [Koch et al., 2018, Turner et al., 2021], reinforcement learning with policy gradient estimation [Malik et al., 2020, Mania et al., 2018], and simulation-based optimization where only function evaluations are accessible [Conn et al., 2009].
@@
Within this setting, two distinct algorithmic paradigms have emerged. The first, often called gradient-approximation methods, constructs surrogate gradient estimates using finite differences and then mimics first-order updates [Nesterov and Spokoiny, 2017, Ghadimi and Lan, 2013, Liu et al., 2018]. Although these approaches inherit convergence guarantees from gradient-based optimization, they introduce discretization bias and require careful tuning of smoothing parameters, which can significantly affect performance in practice.
@@
In contrast, the second paradigm, known as direct-search, operates purely through comparisons of function values at a set of trial points, and updates to the best point without ever forming a gradient estimate [Bergou et al., 2020, Golovin et al., 2020]. This approach is simple to implement and often more robust to noise in function evaluations.
@@
The Stochastic Three Points (STP) method of  [Bergou et al., 2020] is one of the simplest and most efficient forms of direct-search methods, requiring only three function evaluations per iteration and incorporating random search directions (this is where random search comes from). It evaluates the objective at the current iterate and at two origin-symmetric perturbations along a random direction, then selects the best of the three. In the deterministic setting, STP achieves optimal complexity bounds: $O(d/\varepsilon^{2})$ in the smooth non-convex case [Bergou et al., 2020], matching the best-known rate achieved by gradient estimation methods, and $O(d/\varepsilon)$ in the smooth convex case [El Bakkali and Saadi, 2025], matching the best known complexity bound among gradient estimation methods that do not use momentum.
@@
Despite their appeal, STP and random search methods have been analyzed almost exclusively in the deterministic setting, where access to the full objective function is assumed. However, this assumption is often unrealistic in large-scale machine learning, where the objective typically involves a finite sum over millions of data points. Large-scale learning tasks such as deep reinforcement learning or empirical risk minimization require stochastic algorithms. While stochastic gradient estimation methods are well studied, stochastic random search methods remain poorly understood. Yet, their simplicity and freedom from discretization bias make them an attractive alternative to gradient-based approaches. This gap in the theory motivates the development of stochastic variants of random search that can operate effectively with noisy or subsampled objective estimates. Providing sharp guarantees for such methods is therefore both theoretically and practically important. To address this limitation, in [Boucherouite et al., 2024], they proposed Minibatch STP (MiSTP), which replaces the true objective with a minibatch estimate during the three-point comparison of standard STP. Their analysis, however, yields a suboptimal $O(d^{3}/\varepsilon^{6})$ complexity under strong individual smoothness assumptions, significantly worse than the $O(d/\varepsilon^{4})$ rate achieved by gradient-estimation methods like RSGF [Ghadimi and Lan, 2013]. The goal of this work is to close this gap.
@@ 2 Related Work
Direct search and STP. Classical direct-search methods (see [Conn et al., 2009, Vicente, 2013]) rely on deterministic search directions and tend to scale poorly with the dimensionality. In contrast, the STP algorithm achieves improved complexity bounds: $\mathcal{O}(d/\epsilon^{2})$ in the smooth nonconvex setting [Bergou et al., 2020], and $\mathcal{O}(d/\epsilon)$ in the smooth convex setting [El Bakkali and Saadi, 2025]. These results improve upon the $\mathcal{O}(d^{2}/\epsilon^{2})$ and $\mathcal{O}(d^{2}/\epsilon)$ complexities obtained for deterministic direct-search methods [Vicente, 2013, Konečnỳ and Richtárik, 2014]. Follow-up work extended STP with importance sampling [Bibi et al., 2020] and momentum [Gorbunov et al., 2020], but these analyses were conducted only for deterministic objective functions.
STP in the stochastic setting. To the best of our knowledge, the only prior attempt to analyze STP for stochastic objectives is due to Boucherouite et al. [Boucherouite et al., 2024]. Their minibatch STP method achieves a complexity bound of order $\mathcal{O}(d^{3}/\varepsilon^{6})$ under strong individual smoothness assumptions, requiring each stochastic component function $f_{\xi}$ to be smooth. This rate is substantially worse than the $\mathcal{O}(d/\varepsilon^{4})$ complexity known for gradient-estimation-based methods such as RSGF [Ghadimi and Lan, 2013], which approximate gradients using finite differences.
Gradient-estimation methods. Zeroth-order stochastic methods based on gradient approximation (e.g., two-point estimators [Ghadimi and Lan, 2013], ZO-SVRG [Liu et al., 2018]) achieve a complexity bound of ${\mathcal{O}}(d/\varepsilon^{4})$ under individual smoothness assumption. However, these methods introduce discretization bias and require careful tuning of smoothing parameters, which can significantly impact their empirical performance. In contrast, random search methods avoid discretization errors altogether but have historically lacked comparable theoretical guarantees in the stochastic setting.
@@
  • Average smoothness. Under the weaker assumption that only the average function $f$ is $(L_{0},L_{1})$-smooth, we show that Random Search recovers the ${\mathcal{O}}(d^{3}/\varepsilon^{6})$ rate previously obtained in [Boucherouite et al., 2024], but under strictly milder requirements.
  • Finite-sum variance reduction. For finite-sum objectives, we demonstrate that variance reduction can be realized without the memory overhead of storing snapshots as in ZO-SVRG [Liu et al., 2018]. We establish a complexity of ${\mathcal{O}}\!\left(\min\!\Big\{\tfrac{d^{4/3}n^{2/3}}{\varepsilon^{8/3}},\tfrac{dn}{\varepsilon^{2}}\Big\}\right)$, which improves upon the deterministic rate whenever $n\gg d/\varepsilon^{2}$.
  • Helper or human feedback. We extend Random Search to settings where only inexact comparison feedback is available, such as A/B testing [Kohavi et al., 2009] or RLHF [Christiano et al., 2017, Ouyang et al., 2022]. We prove that convergence holds up to an accuracy floor of order ${\mathcal{O}}(\sqrt{d\delta})$, where $\delta$ quantifies the inexactness of the helper or human feedback.
@@ 3.1 Smoothness assumption
We adopt the following notion of smoothness, introduced in [Zhang et al., 2020].
@@
This condition generalizes standard Lipschitz smoothness by allowing the gradient Lipschitz constant to grow proportionally with the gradient norm itself. It has been used to justify adaptive methods such as gradient clipping [Zhang et al., 2020].
@@ 3.2 Random directions
At each iteration the algorithm samples a random direction ${\bm{s}}_{t}\sim\mathcal{D}$. We require the following assumption, adapted from [Bergou et al., 2020, Boucherouite et al., 2024].
@@ 3.3 Algorithm
Stochastic Three Points (STP). STP is a classical derivative-free randomized search method introduced in [Bergou et al., 2020]. At iteration $t$, given an iterate ${\bm{x}}_{t}\in\mathbb{R}^{d}$, a stepsize $\alpha_{t}>0$, and a search direction distribution $\mathcal{D}$ on $\mathbb{R}^{d}$, the method draws ${\bm{s}}_{t}\sim\mathcal{D}$ and forms the two symmetric trial points ${\bm{x}}_{t}^{+}={\bm{x}}_{t}+\alpha_{t}{\bm{s}}_{t},\,\,{\bm{x}}_{t}^{-}={\bm{x}}_{t}-\alpha_{t}{\bm{s}}_{t},$ then evaluates the objective at the three points $\{{\bm{x}}_{t}^{-},\,{\bm{x}}_{t},\,{\bm{x}}_{t}^{+}\}$, and updates according to the best-of-three rule:
@@
Minibatch STP (MiSTP). In the finite-sum setting, MiSTP [Boucherouite et al., 2024] adapts STP by replacing the full objective $f$ with a single empirical objective $f_{\mathcal{C}_{t}}$, computed from a minibatch $\mathcal{C}_{t}$, while retaining the three-point comparison: ${\bm{x}}_{t+1}\in\arg\min_{{\bm{x}}\in\{{\bm{x}}_{t}^{-},\,{\bm{x}}_{t},\,{\bm{x}}_{t}^{+}\}}f_{\mathcal{C}_{t}}({\bm{x}}).$ Because the accept/reject decision is made using $f_{\mathcal{C}_{t}}$, the selected step may worsen the true objective $f$, and thus including the current point ${\bm{x}}_{t}$ does not ensure anymore the monotonic improvement of the true objective as it is the case for classical STP. We drop this baseline evaluation at ${\bm{x}}_{t}$ and adopt a two–point comparison:
@@ 4.1 Case I: Weak average smoothness
We first assume that only the mean function $f$ satisfies $(L_{0},L_{1})$-smoothness (Assumption 1). To control the stochastic error, we assume a bounded variance of the component functions (cf. [Boucherouite et al., 2024]):
@@ 4.2 Case II: Standard sample smoothness
This is the standard assumption in the analysis of stochastic zeroth-order methods (see, e.g., [Ghadimi and Lan, 2013], which corresponds to the special case $L_{1}=0$). Our formulation is slightly more general.
@@ 4.3 Significance of the results
In this regime we match the complexity $\tilde{{\mathcal{O}}}(d^{3}\sigma_{0}^{2}/\varepsilon^{6})$ of [Boucherouite et al., 2024], but under a much weaker assumption: only the average function $f$ is required to be $(L_{0},L_{1})$-smooth, whereas they assumed each $f_{\xi}$ is globally $L$-smooth. Moreover, our analysis shows that the dominant term depends only on $L_{0}$, which can be much smaller than $L=L_{0}+GL_{1}$ where $G$ bounds $\|\nabla f\|$.
@@
Under the standard assumption that each $f_{\xi}$ is $(L_{0},L_{1})$-smooth with respect to $\nabla f$ (as in [Ghadimi and Lan, 2013], $L_{1}=0$), we obtain the improved complexity $\tilde{{\mathcal{O}}}(d\sigma_{1}^{2}/\varepsilon^{4})$. Thus compared to the weak case, both the dependence on dimension and on accuracy improve. This shows that random search achieves the same scaling as gradient estimation methods, while avoiding discretization bias and remaining conceptually simpler.
@@ 6 Learning with Helper or Human Feedback
  • A/B testing: where we cannot measure the exact loss function, but can observe user preferences between two versions of a product or service [Kohavi et al., 2009].
  • Reinforcement learning with human feedback (RLHF): where the reward signal is provided via human preference comparisons rather than explicit numerical scores [Christiano et al., 2017, Ouyang et al., 2022].
@@
that is, the helper provides a $\delta$-accurate approximation of the true difference in function values. We call this property $\delta$-inexact feedback. This is similar in spirit to the helper framework introduced in [Chayti and Karimireddy, 2024] for first-order methods and [Chayti et al., 2024] for second-order methods.
@@ 7 On the Limitations of Momentum for Random Search
A natural question is whether the variance-reduction ideas developed for first-order stochastic optimization can be transferred to Random Search. In the first-order literature, several approaches exploit momentum to mitigate variance, including Heavy Ball [Polyak, 1964], momentum-based variance reduction such as STORM [Cutkosky and Orabona, 2019], and implicit gradient transport [Arnold et al., 2019]. These methods help precisely when gradient noise is high and batches are small.
@@ Conclusion.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [No] we are open to it even though we don’t see a need since our experiments are easy.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes] in the appendix
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Yes/No/Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Appendix A Preliminaries
Let $(Z_{i})_{i=1}^{b}$ be i.i.d. real-valued random variables with $\mathbb{E}[Z_{1}]=0$ and $\mathrm{Var}(Z_{1})\leq\sigma^{2}<\infty$. Define the minibatch average $\overline{Z}:=\tfrac{1}{b}\sum_{i=1}^{b}Z_{i}$. Then
$\mathbb{E}[\overline{Z}]=0\qquad\text{and}\qquad\mathrm{Var}(\overline{Z})=\frac{\mathrm{Var}(Z_{1})}{b}\;\leq\;\frac{\sigma^{2}}{b}.$
@@
Let $\phi:\mathbb{R}\to\mathbb{R}$ be convex and let $X$ be an integrable real-valued random variable. Then
$\phi\big(\mathbb{E}[X]\big)\;\leq\;\mathbb{E}\big[\phi(X)\big].$
In particular, for $\phi(x)=|x|$ we have $|\mathbb{E}[X]|\leq\mathbb{E}[|X|]$.
@@
Consider a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfying the following property for all ${\bm{x}},{\bm{x}}^{\prime}\in\mathbb{R}^{d}$:
$\big\|\nabla f({\bm{x}})-\nabla f({\bm{x}}^{\prime})\big\|\leq\min\big({\mathcal{L}}({\bm{x}}),{\mathcal{L}}({\bm{x}}^{\prime})\big)\|{\bm{x}}-{\bm{x}}^{\prime}\|,$
for some function ${\mathcal{L}}:\mathbb{R}^{d}\to\mathbb{R}^{+}$. Then, the function value is bounded by the quadratic approximation at $\bm{x}$} as follows:
$\big|f({\bm{x}}^{\prime})-f({\bm{x}})-\nabla f({\bm{x}})^{\top}({\bm{x}}^{\prime}-{\bm{x}})\big|\leq\dfrac{{\mathcal{L}}({\bm{x}})}{2}\|{\bm{x}}^{\prime}-{\bm{x}}\|^{2}.$

### v2

@@ 1 Introduction
Zeroth-order (derivative-free) optimization refers to a class of algorithms that minimize an objective function using only function evaluations. These methods are crucial in modern machine learning applications, where gradients are either inaccessible, unreliable, or too costly to obtain. Such scenarios arise in black-box adversarial attacks [chen2017zoo, ughi2022empirical], hyperparameter tuning [koch2018autotune, turner2021bayesian], reinforcement learning with policy gradient estimation [malik2020derivative, mania2018simple], and simulation-based optimization where only function evaluations are accessible [conn2009introduction].
@@
Within this setting, two distinct algorithmic paradigms have emerged. The first, often called gradient-approximation methods, constructs surrogate gradient estimates using finite differences and then mimics first-order updates [nesterov2017random, ghadimi2013stochastic, liu2018zeroth]. Although these approaches inherit convergence guarantees from gradient-based optimization, they introduce discretization bias and require careful tuning of smoothing parameters, which can significantly affect performance in practice.
@@
In contrast, the second paradigm, known as direct-search, operates purely through comparisons of function values at a set of trial points, and updates to the best point without ever forming a gradient estimate [bergou2020stochastic, golovin2020gradientless]. This approach is simple to implement and often more robust to noise in function evaluations.
@@
The Stochastic Three Points (STP) method of  [bergou2020stochastic] is one of the simplest and most efficient forms of direct-search methods, requiring only three function evaluations per iteration and incorporating random search directions (this is where random search comes from). It evaluates the objective at the current iterate and at two origin-symmetric perturbations along a random direction, then selects the best of the three. In the deterministic setting, STP achieves optimal complexity bounds: $O(d/\varepsilon^{2})$ in the smooth non-convex case [bergou2020stochastic], matching the best-known rate achieved by gradient estimation methods, and $O(d/\varepsilon)$ in the smooth convex case [kadi2025on], matching the best known complexity bound among gradient estimation methods that do not use momentum.
@@
Despite their appeal, STP and random search methods have been analyzed almost exclusively in the deterministic setting, where access to the full objective function is assumed. However, this assumption is often unrealistic in large-scale machine learning, where the objective typically involves a finite sum over millions of data points. Large-scale learning tasks such as deep reinforcement learning or empirical risk minimization require stochastic algorithms. While stochastic gradient estimation methods are well studied, stochastic random search methods remain poorly understood. Yet, their simplicity and freedom from discretization bias make them an attractive alternative to gradient-based approaches. This gap in the theory motivates the development of stochastic variants of random search that can operate effectively with noisy or subsampled objective estimates. Providing sharp guarantees for such methods is therefore both theoretically and practically important. To address this limitation, in [boucherouite2024mistp], they proposed Minibatch STP (MiSTP), which replaces the true objective with a minibatch estimate during the three-point comparison of standard STP. Their analysis, however, yields a suboptimal $O(d^{3}/\varepsilon^{6})$ complexity under strong individual smoothness assumptions, significantly worse than the $O(d/\varepsilon^{4})$ rate achieved by gradient-estimation methods like RSGF [ghadimi2013stochastic]. The goal of this work is to close this gap.
@@ 2 Related Work
Direct search and STP. Classical direct-search methods (see [conn2009introduction, vicente2013worst]) rely on deterministic search directions and tend to scale poorly with the dimensionality. In contrast, the STP algorithm achieves improved complexity bounds: $\mathcal{O}(d/\epsilon^{2})$ in the smooth nonconvex setting [bergou2020stochastic], and $\mathcal{O}(d/\epsilon)$ in the smooth convex setting [kadi2025on]. These results improve upon the $\mathcal{O}(d^{2}/\epsilon^{2})$ and $\mathcal{O}(d^{2}/\epsilon)$ complexities obtained for deterministic direct-search methods [vicente2013worst, konevcny2014simple]. Follow-up work extended STP with importance sampling [bibi2020stochastic] and momentum [gorbunov2020stochastic], but these analyses were conducted only for deterministic objective functions.
STP in the stochastic setting. To the best of our knowledge, the only prior attempt to analyze STP for stochastic objectives is due to Boucherouite et al. [boucherouite2024mistp]. Their minibatch STP method achieves a complexity bound of order $\mathcal{O}(d^{3}/\varepsilon^{6})$ under strong individual smoothness assumptions, requiring each stochastic component function $f_{\xi}$ to be smooth. This rate is substantially worse than the $\mathcal{O}(d/\varepsilon^{4})$ complexity known for gradient-estimation-based methods such as RSGF [ghadimi2013stochastic], which approximate gradients using finite differences.
Gradient-estimation methods. Zeroth-order stochastic methods based on gradient approximation (e.g., two-point estimators [ghadimi2013stochastic], ZO-SVRG [liu2018zeroth]) achieve a complexity bound of ${\mathcal{O}}(d/\varepsilon^{4})$ under individual smoothness assumption. However, these methods introduce discretization bias and require careful tuning of smoothing parameters, which can significantly impact their empirical performance. In contrast, random search methods avoid discretization errors altogether but have historically lacked comparable theoretical guarantees in the stochastic setting.
@@
  • Average smoothness. Under the weaker assumption that only the average function $f$ is $(L_{0},L_{1})$-smooth, we show that Random Search recovers the ${\mathcal{O}}(d^{3}/\varepsilon^{6})$ rate previously obtained in [boucherouite2024mistp], but under strictly milder requirements.
  • Finite-sum variance reduction. For finite-sum objectives, we demonstrate that variance reduction can be realized without the memory overhead of storing snapshots as in ZO-SVRG [liu2018zeroth]. We establish a complexity of ${\mathcal{O}}\!\left(\min\!\Big\{\tfrac{d^{4/3}n^{2/3}}{\varepsilon^{8/3}},\tfrac{dn}{\varepsilon^{2}}\Big\}\right)$, which improves upon the deterministic rate whenever $n\gg d/\varepsilon^{2}$.
  • Helper or human feedback. We extend Random Search to settings where only inexact comparison feedback is available, such as A/B testing [kohavi2009controlled] or RLHF [christiano2017deep, ouyang2022training]. We prove that convergence holds up to an accuracy floor of order ${\mathcal{O}}(\sqrt{d\delta})$, where $\delta$ quantifies the inexactness of the helper or human feedback.
@@ 3.1 Smoothness assumption
We adopt the following notion of smoothness, introduced in [zhang2020gradientclippingacceleratestraining].
@@
This condition generalizes standard Lipschitz smoothness by allowing the gradient Lipschitz constant to grow proportionally with the gradient norm itself. It has been used to justify adaptive methods such as gradient clipping [zhang2020gradientclippingacceleratestraining].
@@ 3.2 Random directions
At each iteration the algorithm samples a random direction ${\bm{s}}_{t}\sim\mathcal{D}$. We require the following assumption, adapted from [bergou2020stochastic, boucherouite2024mistp].
@@ 3.3 Algorithm
Stochastic Three Points (STP). STP is a classical derivative-free randomized search method introduced in [bergou2020stochastic]. At iteration $t$, given an iterate ${\bm{x}}_{t}\in\mathbb{R}^{d}$, a stepsize $\alpha_{t}>0$, and a search direction distribution $\mathcal{D}$ on $\mathbb{R}^{d}$, the method draws ${\bm{s}}_{t}\sim\mathcal{D}$ and forms the two symmetric trial points ${\bm{x}}_{t}^{+}={\bm{x}}_{t}+\alpha_{t}{\bm{s}}_{t},\,\,{\bm{x}}_{t}^{-}={\bm{x}}_{t}-\alpha_{t}{\bm{s}}_{t},$ then evaluates the objective at the three points $\{{\bm{x}}_{t}^{-},\,{\bm{x}}_{t},\,{\bm{x}}_{t}^{+}\}$, and updates according to the best-of-three rule:
@@
Minibatch STP (MiSTP). In the finite-sum setting, MiSTP [boucherouite2024mistp] adapts STP by replacing the full objective $f$ with a single empirical objective $f_{\mathcal{C}_{t}}$, computed from a minibatch $\mathcal{C}_{t}$, while retaining the three-point comparison: ${\bm{x}}_{t+1}\in\arg\min_{{\bm{x}}\in\{{\bm{x}}_{t}^{-},\,{\bm{x}}_{t},\,{\bm{x}}_{t}^{+}\}}f_{\mathcal{C}_{t}}({\bm{x}}).$ Because the accept/reject decision is made using $f_{\mathcal{C}_{t}}$, the selected step may worsen the true objective $f$, and thus including the current point ${\bm{x}}_{t}$ does not ensure anymore the monotonic improvement of the true objective as it is the case for classical STP. We drop this baseline evaluation at ${\bm{x}}_{t}$ and adopt a two–point comparison:
@@ 4.1 Case I: Weak average smoothness
We first assume that only the mean function $f$ satisfies $(L_{0},L_{1})$-smoothness (Assumption 1). To control the stochastic error, we assume a bounded variance of the component functions (cf. [boucherouite2024mistp]):
@@ 4.2 Case II: Standard sample smoothness
This is the standard assumption in the analysis of stochastic zeroth-order methods (see, e.g., [ghadimi2013stochastic], which corresponds to the special case $L_{1}=0$). Our formulation is slightly more general.
@@ 4.3 Significance of the results
In this regime we match the complexity $\tilde{{\mathcal{O}}}(d^{3}\sigma_{0}^{2}/\varepsilon^{6})$ of [boucherouite2024mistp], but under a much weaker assumption: only the average function $f$ is required to be $(L_{0},L_{1})$-smooth, whereas they assumed each $f_{\xi}$ is globally $L$-smooth. Moreover, our analysis shows that the dominant term depends only on $L_{0}$, which can be much smaller than $L=L_{0}+GL_{1}$ where $G$ bounds $\|\nabla f\|$.
@@
Under the standard assumption that each $f_{\xi}$ is $(L_{0},L_{1})$-smooth with respect to $\nabla f$ (as in [ghadimi2013stochastic], $L_{1}=0$), we obtain the improved complexity $\tilde{{\mathcal{O}}}(d\sigma_{1}^{2}/\varepsilon^{4})$. Thus compared to the weak case, both the dependence on dimension and on accuracy improve. This shows that random search achieves the same scaling as gradient estimation methods, while avoiding discretization bias and remaining conceptually simpler.
@@ 6 Learning with Helper or Human Feedback
  • A/B testing: where we cannot measure the exact loss function, but can observe user preferences between two versions of a product or service [kohavi2009controlled].
  • Reinforcement learning with human feedback (RLHF): where the reward signal is provided via human preference comparisons rather than explicit numerical scores [christiano2017deep, ouyang2022training].
@@
that is, the helper provides a $\delta$-accurate approximation of the true difference in function values. We call this property $\delta$-inexact feedback. This is similar in spirit to the helper framework introduced in [chayti2024optimization] for first-order methods and [chayti2024unified] for second-order methods.
@@ 7 On the Limitations of Momentum for Random Search
A natural question is whether the variance-reduction ideas developed for first-order stochastic optimization can be transferred to Random Search. In the first-order literature, several approaches exploit momentum to mitigate variance, including Heavy Ball [polyak1964some], momentum-based variance reduction such as STORM [cutkosky2019momentum], and implicit gradient transport [arnold2019reducing]. These methods help precisely when gradient noise is high and batches are small.
@@ Appendix A Preliminaries
Let $(Z_{i})_{i=1}^{b}$ be i.i.d. real-valued random variables with $\mathbb{E}[Z_{1}]=0$ and $\mathrm{Var}(Z_{1})\leq\sigma^{2}<\infty$. Define the minibatch average $\overline{Z}:=\tfrac{1}{b}\sum_{i=1}^{b}Z_{i}$. Then $\mathbb{E}[\overline{Z}]=0\qquad\text{and}\qquad\mathrm{Var}(\overline{Z})=\frac{\mathrm{Var}(Z_{1})}{b}\;\leq\;\frac{\sigma^{2}}{b}.$
@@
Let $\phi:\mathbb{R}\to\mathbb{R}$ be convex and let $X$ be an integrable real-valued random variable. Then $\phi\big(\mathbb{E}[X]\big)\;\leq\;\mathbb{E}\big[\phi(X)\big].$ In particular, for $\phi(x)=|x|$ we have $|\mathbb{E}[X]|\leq\mathbb{E}[|X|]$.
@@
Consider a differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfying the following property for all ${\bm{x}},{\bm{x}}^{\prime}\in\mathbb{R}^{d}$: $\big\|\nabla f({\bm{x}})-\nabla f({\bm{x}}^{\prime})\big\|\leq\min\big({\mathcal{L}}({\bm{x}}),{\mathcal{L}}({\bm{x}}^{\prime})\big)\|{\bm{x}}-{\bm{x}}^{\prime}\|,$ for some function ${\mathcal{L}}:\mathbb{R}^{d}\to\mathbb{R}^{+}$. Then, the function value is bounded by the quadratic approximation at $\bm{x}$} as follows: $\big|f({\bm{x}}^{\prime})-f({\bm{x}})-\nabla f({\bm{x}})^{\top}({\bm{x}}^{\prime}-{\bm{x}})\big|\leq\dfrac{{\mathcal{L}}({\bm{x}})}{2}\|{\bm{x}}^{\prime}-{\bm{x}}\|^{2}.$